# Exploring hidden priors when interpreting gravitational wave and electromagnetic probes of the nuclear equation of state

### v1

@@ III.3 On the maximum neutron star mass and pertinent information that limits it
Figure 13: Impact of maximum nonrotating NS mass: Top panel: Posterior cumulative distribution for maximum nonrotating NS mass: The fraction of EoS in our posterior with $M_{\rm max}$ less than a given threshold, versus that threshold. The blue curve shows the result for our fiducial analysis with the $\Gamma$-spectral EoS with galactic pulsars and the symmetry energy, as shown in Figure 8. The orange curve shows the corresponding result after applying a nominal likelihood designed to disfavor large maximum masses, as shown in the bottom panel of Figure 6. Bottom panel: Marginal likelihood for different EoS, versus $R_{1.4}$. These likelihood evaluations correspond to the orange curve in the top panel.
@@
Despite being quite restricted in form and range, our two headline equation of state families allow for EoS with very large maximum neutron star mass $M_{max}$; see, e.g., Figure 13. The observations of galactic pulsars described above do not identify an upper bound on the neutron star maximum mass, explicitly or implicitly; see, e.g., [55] and references therein. Within our framework, this degeneracy manifests as a broad plateau of the likelihood that covers EoS hyperparameters including such extreme scenarios, as required by this observational constraint; see Figure 6. In our framework these EoS usually exhibit extreme behavior at high density, such as $c_{s}\rightarrow 1$.
@@
Several investigations have proposed theoretically or observationally motivated limits on the EoS which would disfavor these high-$M_{max}$ EoS. In the most extreme approach, Ref. [55] suggests that AT2017gfo’s energetics require a prompt collapse to form a black hole, limiting the neutron star maximum mass to be below $2.16^{+0.17}_{-0.15}$ M. The bottom panel of Figure 6 shows the specific likelihood we adopt to impose this constraint, following previous work [28].
@@
To illustrate the impact of this suggestion, Figure 14 shows updated results from our fiducial NICER analysis, now incorporating this proposed upper limit. As desired, these constraints disfavor large $M_{max}$. That said, taken with our other constraints, our analysis still allows for the possibility that $M_{\rm max}$ is significantly larger than the median estimate proposed by [55] using this prior information alone, and conversely strongly disfavors the smallest maximum masses allowed a priori. As a critical side effect illustrated in Figure 14, these constraints on $M_{max}$ also inform the EoS over its full dynamic range, shifting and narrowing our posterior distributions in $p(\rho)$ and $c_{s}(\rho)$, particularly at densities well above nuclear saturation density.
@@ III.4 Adding the HESS source
Above, we have emphasized EoS inferences using conventional galactic pulsars, explicitly omitting the HESS source J1731 [83] shown in Figure 2. Motivated by previous investigations to assess the impact of this event on EoS inference [152, 153, 154, 155, 156, 157], with Figure 15 we show the results of EoS inference which explicitly include this event, compared with our previous results for galactic pulsars only. Despite the nominal low maximum-likelihood mass and radius deduced for this event, when accounting for the full marginal likelihood we find GW measurements in particular are very consistent with the HESS likelihood. As a result, adding the HESS event produces only incremental changes in the EoS, both as shown here and when adding in GW and other information. In particular, the HESS observation does not seem to require exceptional EoS choices.
Figure 15: Incremental impact of HESS source: Like Figure 8, but adding the HESS source. Except for a modest change at small radius, the overall EoS is only incrementally changed when the HESS source is included, largely at low mass. We also provide a benchmark analysis with the same PSRs and symmetry energy and uses the same $\Gamma$-spectral parameterization, but also includes GW constraint, for these constraints [45].
@@ IV Conclusions
In this proof-of-concept study, we have investigated how several often-obscured sources of prior information propagate into statements about the nuclear equation of state. Most notably, we have demonstrated that hidden astrophysical population assumptions about the source responsible for GW170817/AT2017gfo directly propagate into substantially different conclusions, both about the radius of a fiducial neutron star and more broadly about the nuclear equation of state. In so doing, we have clarified how EoS inference requires and explicitly depends on strong assumptions about each pertinent neutron star population. Using two simple but different EoS parameterizations, we provide a concrete demonstration of how severely conclusions about the nuclear EoS can depend on parameterizations and prior choices, exemplifying conclusions identified in prior work. To concretely illuminate the impact of the prior, our hyperparameter plots indicate marginal EoS likelihoods; in every single analysis, many EoS with the highest marginal likelihood are not within our hyperparameter posterior, being disfavored by our hyperprior. Using the marginal likelihood to separate the impact of EoS priors, we show many EoS in our two model families are both consistent with current pulsar and GW observations but support very large NS masses. Accounting for these hidden systematics, we argue the nuclear equation of state is much less well constrained by extant observations than previously reported, particularly by investigations which stack multiple heterogeneous measurements whose relative systematics are not well understood.
@@
Our conclusions are compatible with other recent reassessments of EoS-inference-adjacent observational and theoretical systematics [158]. Notably, recent re-examination of NICER observations of J0030+4051 with a more sophisticated approach [159] unveils a wider range of possible interpretations than the widely-used unimodal likelihood employed as input in this work.
@@
In this investigation, we have also methodically investigated several proposed sources for external information: galactic pulsar measurements of mass, mass and radius, including the HESS measurement of PSR J1731, GW measurements from GW170817, the nuclear symmetry energy implied by PREX, an upper bound on the NS maximum mass, motivated by energetics of AT2017gfo, and proposed lower bounds on $\tilde{\Lambda}$ also motivated by this event. Using the (range of the) marginal log-likelihoods involved as a diagnostic of information provided by each measurement, we argue that the primary measurements (galactic pulsars, including the HESS source; GW measurements; and the symmetry energy) are remarkably consistent for our model family, particularly when allowing for systematic uncertainty in the interpretation of GW170817 due to hidden astrophysical population assumptions. This consistency suggests observations presently only coarsely constrain the EoS space, leaving limited opportunity for more flexible EoS to arrive at qualitatively different conclusions. Thus, rather than use modest tension between different observations (e.g., fidicual GW versus NICER) as a mechanism to draw narrower conclusions about the NS radius, which requires well-controlled systematics, we instead argue these tensions point to wider posterior conclusions than previously reported. Specifically, taking the union of our credible intervals between the most optimistic and pessimistic scenarios allowing for both choices for EoS family, we argue $R_{1.4}$ is between $11{\rm km}$ and $14{\rm km}$; see Figures 11 and 9. This range is qualitatively consistent with a simple, pessimistic heuristic interpretation of how systematics limit our observations, when analyzed jointly and accounting for systematics: individual measurements’ maximum likelihood estimates bracket the available range. We emphasize that these optimistic and pessimistic conclusions are driven in no small part by assumptions: which events and interpretations are used, with what EoS family and priors.
@@
In our proof-of-concept study, we have intentionally adopted a fixed low-density EoS and thus a largely known NS radius at low mass. By adopting such a strong prior, we guarantee that low-mass NS will have the expected radius and limit the possibility that even extreme choices for our EoS parameters can be in tension with observations. In other words, by fixing the low-density EoS, we increase the chance that the radius $R_{1.4}$ will be consistent with most measurements, leaving other phenomenological coordinates like $M_{max}$ as free (and observationally unbounded) parameters. While in this work for simplicity we have adopted a fixed low-dimensional nuclear equation of state, we recognize that the high precision of future EoS constraints requires a more flexible, powerful framework that incorporates more (perhaps infinitely many) dimensions and allows for more flexible physics like phase transitions.
@@
To carry out our calculations, we have deployed an updated version of the RIFT pipeline capable of efficiently performing hierarchical inference with multiple objects, multiple messengers, and even flexible source populations to infer a joint equation of state. In this work, we built our analysis pipeline using fixed hierarchical inputs: here, approximated NICER likelihoods and high-precision marginal GW likelihoods. In future work we will generalize this pipeline to reassess the hierarchical inputs and request further refinement when needed (e.g., more marginal likelihood evaluations from the GW analysis, to better resolve the low-$\tilde{\Lambda}$ region if needed).
@@
Finally, in an appendix, we outline a simple estimate which estimates how well future GW surveys can constrain the nuclear equation of state. This estimate demonstrates that most individual NS observations will directly inform the nuclear EoS, not just the loudest few. This estimate also demonstrates that, due to the strong trend of tidal deformability with radius, the least massive NS should dominate conclusions about the EoS, excepting only that nature produces them too infrequently.
@@
The most surprising new result in our study is the strong dependence of GW-inferred tidal deformability on astrophysical prior assumptions. In addition to a fiducial analysis of GW170817, we also analyze that event using two well-motivated binary NS population models. In one model, we require GW170817 be more similar to the expected binary neutron star population, based on theoretical models for their formation: a comparable mass ratio, and with spins preferentially positively aligned with the orbital angular momentum. In the other model, designed to produce (much) larger amounts of radioactive ejecta, we require a highly asymmetric binary (potentially rapidly spinning). These two prior assumptions notably change the conclusions derived from GW170817 and hence the conclusions derived when combining this event with other analyses. Though the latter of our two scenarios seems in tension with the long-lived galactic BNS population, we note too that a rapidly-merging subpopulation could have different population properties, and offers an attractive resolution to the long-term challenge of forming r-process elements. Conversely, if we adopt a conventional galactic BNS origin for 170817 and incorporate all available measurements, we conclude $R_{1.4}\simeq 11.8{\rm km}-12.9{\rm km}$.
@@
For simplicity, in this proof-of-concept analysis we have omitted several pertinent areas of investigations. We have notably adopted a single GW waveform model, eliding the effect of waveform modeling systematics. Too, given substantial modeling systematics, we have also only adopted toy models for the information provided from kilonova observations which limit the possible BNS mass, spin, and EoS. We have taken existing all fiducial galactic pulsar measurements at face value; reanalyses of these observations may point to larger systematics and broader posteriors than adoped here [159]. Finally, we have adopted an intentionally simple EoS model family and prior, omitting often-discussed constraints available at low density (e.g., from chiral effective field theory and from state-of-the-art crust models) and at high density (e.g, that high-density EoS should asymptotically attain the perturbative-QCD limit of $c_{s}\rightarrow 1/\sqrt{3}$ approaching from below [160, 161]). Our choice of these simplifications enables us to focus exclusively on the often strong and hidden impact that hidden astrophysical priors produce.
@@
The multifactor iterative inference algorithm employed in this work is available as the “HyperPipe” component of the RIFT python software package, available at https://pypi.org/project/RIFT/. Instructions for this approach are available at [162].
@@ Appendix A NICER data and Gaussian approximation
The NICER data releases [163, 164, 165, 166] provide weighted posterior samples $(x_{k},w_{k})$ derived from comparing their pulse models and neutron star mass-radius combinations with observed X-ray photons. We fit a conventional two-dimensional normal distribution to these weighted samples; see, e.g., the discussion in  For J0030, which lacks any hard edges introduced by causality constraints in the $(M,R)$ plane, we use the weighted empirical mean $\bar{x}=\sum_{k}w_{k}x_{k}/\sum_{k}w_{k}$ and weighted empirical covariance matrix to infer the parameters of the best-fitting gaussian. More generally, however, we optimize the mean $\mu$ and covariance matrix $\Sigma$ of a standard normal distribution, according to
$\ell(\mu,\Sigma)=\sum_{k}w_{k}\ln p(x_{k}|\mu,\Sigma)$ (13)
[In the limit of many weighted samples constructed by importance sampling, with $w_{k}=L(x_{k})$, the quantity $\ell$ will approach the KL divergence between the (truncated) Gaussian and the full posterior.] As a check on our results and their stability, we repeat our analysis with fair draws from the weighted samples, using the expression above with $w_{k}=1$. The best fitting normal distribution in $(R/{\rm km},M/M_{\odot})$ has mean $[13.345373961.44968231]$ and covariance matrix
$\begin{bmatrix}1.52687831&0.18343073\\ 0.18343073&0.02778826\end{bmatrix}$

### v2

@@ III.3 On the maximum neutron star mass and pertinent information that limits it
Figure 13: Impact of maximum nonrotating NS mass: Top panel: Posterior cumulative distribution for maximum nonrotating NS mass: The fraction of EoS in our posterior with $M_{\rm max}$ less than a given threshold, versus that threshold. The blue curve shows the result for our fiducial analysis with the $\Gamma$-spectral EoS with galactic pulsars and the symmetry energy, as shown in Figure 5. The orange curve shows the corresponding result after applying a nominal likelihood designed to disfavor large maximum masses, as shown in the bottom panel of Figure 4. Bottom panel: Marginal likelihood for different EoS, versus Mmax. These likelihood evaluations correspond to the orange curve in the top panel.
@@
Despite being quite restricted in form and range, our two headline equation of state families allow for EoS with very large maximum neutron star mass $M_{max}$; see, e.g., Figure 13. The observations of galactic pulsars described above do not identify an upper bound on the neutron star maximum mass, explicitly or implicitly; see, e.g., [55] and references therein. Within our framework, this degeneracy manifests as a broad plateau of the likelihood that covers EoS hyperparameters including such extreme scenarios, as required by this observational constraint; see Figure 4. In our framework these EoS usually exhibit extreme behavior at high density such as their $c_{s}\rightarrow 1$.
@@
Several investigations have proposed theoretically or observationally motivated limits on the EoS which would disfavor these high-$M_{max}$ EoS. In the most extreme approach, Ref. [55] suggests that AT2017gfo’s energetics require a prompt collapse to form a black hole, limiting the neutron star maximum mass to be below $2.16^{+0.17}_{-0.15}~{}{\rm M_{\odot}}$. The bottom panel of Figure 4 shows the specific likelihood we adopt to impose this constraint, following previous work [28].
@@
To illustrate the impact of this suggestion, Figure 14 shows updated results from our fiducial NICER analysis, now incorporating this proposed upper limit. As desired, these constraints disfavor large $M_{max}$. That said, taken with our other constraints, our analysis still allows for the possibility that $M_{\rm max}$ is larger than the median estimate proposed by [55] using this prior information alone, and conversely still disfavors the smallest maximum masses allowed a priori due to the strongly implied minimums on $M_{max}$ through observed NSs.
We notice that there is a shift of the M-R curve at large masses towards smaller radii which indicates that favored EoSs have a M-R that bends towards the causality limit boundary at the top left, and hence implies restrictions on EoSs. Note that these boundaries are obtained by evaluating the 90% percentile credible intervals at each mass, and hence will always have a posterior at high masses even though the EoSs with highest likelihoods do not reach such higher mass values.
As a critical side effect, these constraints on $M_{max}$ also inform the EoS over its full dynamic range, shifting and narrowing our posterior distributions in $p(\rho)$ and $c_{s}(\rho)$, particularly at densities well above nuclear saturation density. Further, the soundspeed distribution narrows to higher speeds at high densities in order to compensate for the reduced stiffness at intermediate densities with lower sound speeds and lower pressures to provide with larger pressures that produce stable cores for the substantially compactified star. This results in an M-R that reach Masses of up to $\sim 2~{}{\rm M_{\odot}}$ but does not surpass it substantially.
@@ III.4 Adding the HESS source
In the main text, we have emphasized EoS inferences using conventional galactic pulsars, explicitly omitting the HESS source J1731 [83] shown in Figure 2. The inferred radius of this source is subject to strong systematic uncertainties [155]. Motivated by previous investigations to assess the impact of this event on EoS inference [156, 157, 158, 159, 160, 161], with Figure 15 we show the results of EoS inference which explicitly include this event, compared with our previous results for galactic pulsars only. Despite the nominal low maximum-likelihood mass and radius deduced for this event, when accounting for the full marginal likelihood we find GW measurements in particular are very consistent with the HESS likelihood. As a result, adding the HESS event produces only incremental changes in the EoS, both as shown here and when adding in GW and other information. In particular, the HESS observation does not seem to require exceptional EoS choices.
Figure 15: Incremental impact of HESS source: Like Figure 5, but adding the HESS source. Except for a modest change at small radius, the overall EoS is only incrementally changed when the HESS source is included, largely at low mass. A benchmark analysis with the same PSRs and symmetry energy and uses the same $\Gamma$-spectral parameterization (also includes GW constraint) is also provided [45].
@@ IV Conclusions
In this proof-of-concept study, we have investigated how several often-obscured sources of prior information propagate into statements about the nuclear equation of state. Most notably, we have demonstrated that hidden astrophysical population assumptions about the source responsible for GW170817/AT2017gfo directly propagate into substantially different conclusions, both about the radius of a fiducial neutron star and more broadly about the nuclear equation of state. In doing so, we have clarified how EoS inference requires and explicitly depends on strong assumptions about each pertinent neutron star population. Using two EoS parameterizations, we show how severely conclusions about the nuclear EoS can depend on parameterizations and prior choices. To concretely illuminate the impact of the prior, our hyperparameter plots indicate marginal EoS likelihoods; in every analysis, many EoSs with the highest marginal likelihood are not within our hyperparameter posterior, being disfavored by the hyperparameter prior (i.e. the priors on the two spectral parameterizations). Using the marginal likelihood to separate the impact of EoS priors, we show many EoS in our two model families are both consistent with current pulsar and GW observations but support very large NS masses. Accounting for these hidden systematics, we argue the nuclear equation of state is much less well constrained by extant observations than previously reported, particularly by investigations which stack multiple heterogeneous measurements whose relative systematics are not well understood.
The $\Upsilon$-spectral model for EoS attains a higher marginal likelihood than $\Gamma$-spectral model does, implying that specific EOS samples produced by the $\Upsilon$-spectral parameterization fits the observational constraint better than $\Gamma$-spectral. This is because $\Upsilon$-spectral is able to produce EoSs with rapid variation in pressure with density increase as can be seen in Fig. 6. Other parameterizations such as $\Gamma$-spectral and even piecewise-polytropic do not form EoSs that can show rapid shift in slope due to the specific functional forms, and hence produce EoSs with a weaker fit. The implication of $\Upsilon$-spectral parameterization is that it infers a strong phase-transition-like feature at $\sim 10^{14.5}$ g cm-3.
@@
Our conclusions are compatible with other recent reassessments of EoS-inference-adjacent observational and theoretical systematics [162]. Notably, recent re-examination of NICER observations of J0030+4051 [163] unveils a wider range of possible interpretations than the widely-used unimodal likelihood employed as input in this work. Our conclusions are also compatible with a recent galactic pulsar observations of J0437, as summarized in an approximate analysis in Appendix E [142].
@@
We have methodically investigated the impact on EoS inference by several external sources of information: galactic pulsar measurements of mass, mass and radius, including the HESS measurement of PSR J1731, GW measurements from GW170817, the nuclear symmetry energy implied by PREX, an upper bound on the NS maximum mass, motivated by energetics of AT2017gfo, and discussed proposed lower bounds on $\tilde{\Lambda}$ also motivated by this event. Using the (range of the) marginal log-likelihoods involved as a diagnostic of information provided by each measurement we argue that the primary measurements (galactic pulsars, including the HESS source; GW measurements; and the symmetry energy) are remarkably consistent for our model family, particularly when allowing for systematic uncertainty in the interpretation of GW170817 due to hidden astrophysical population assumptions. This consistency suggests observations presently only coarsely constrain the EoS space, leaving limited opportunity for more flexible EoS to arrive at qualitatively different conclusions. Thus, rather than use modest tension between different observations (e.g., fidicual GW versus NICER) as a mechanism to draw narrower conclusions about the NS radius, which requires well-controlled systematics, we instead argue these tensions point to wider posterior conclusions than previously reported. Specifically, taking the union of our credible intervals between the most optimistic and pessimistic scenarios allowing for both choices for EoS family, we argue $R_{1.4}$ is between $11~{}{\rm km}$ and $14~{}{\rm km}$; see Figures 10 and 6. This range is qualitatively consistent with a simple, pessimistic heuristic interpretation of how systematics limit our observations, when analyzed jointly and accounting for systematics: individual measurements’ maximum likelihood estimates bracket the available range. We emphasize that these optimistic and pessimistic conclusions are driven in no small part by assumptions: which events and interpretations are used, with what EoS family and priors.
@@
In our proof-of-concept study, we have intentionally adopted a fixed low-density EoS and thus a largely known NS radius at low mass. By adopting such a strong prior, we guarantee that low-mass NS will have the expected radius and limit the possibility that even extreme choices for EoS parameters can be in tension with observations. In other words, by fixing the low-density EoS, we increase the chance that the radius $R_{1.4}$ will be consistent with most measurements, leaving other phenomenological coordinates like $M_{max}$ as free (and observationally unbounded) parameters. While in this work for simplicity we have adopted a fixed low-dimensional nuclear equation of state, we recognize that the high precision of future EoS constraints requires a more flexible, powerful framework that incorporates more dimensions and allows for more flexible physics like phase transitions.
@@
To carry out our calculations, we have deployed an updated version of the RIFT pipeline capable of efficiently performing hierarchical inference with multiple objects, multiple messengers, and even flexible source populations to infer a joint equation of state. We built our analysis pipeline using fixed hierarchical inputs: here, approximated NICER likelihoods and high-precision marginal GW likelihoods. In future work we will generalize this pipeline to reassess the hierarchical inputs and request further refinement when needed (e.g., more marginal likelihood evaluations from the GW analysis, to better resolve the low-$\tilde{\Lambda}$ region if needed).
@@
In Appendix B, we outline a simple estimate of how well future GW surveys can constrain the nuclear equation of state. This estimate demonstrates that most individual NS observations will directly inform the nuclear EoS, not just the loudest few. This estimate also demonstrates that, due to the strong trend of tidal deformability with radius, the least massive NS should dominate conclusions about the EoS, excepting only that nature produces them too infrequently. Additionally, this appendix illustrates how EOS constraints will depend on two key assumptions held fixed in the main text: the assumption of a fixed, uniform NS mass distribution on the one hand, and on the number and distance distribution on the other.
@@
The most surprising new result in our study is the strong dependence of GW-inferred tidal deformability on astrophysical prior assumptions. In addition to a fiducial analysis of GW170817, we also analyze this event using two well-motivated binary NS population models. In one model, we require GW170817 be more similar to the expected binary neutron star population, based on theoretical models for their formation: a comparable mass ratio, and with spins preferentially positively aligned with the orbital angular momentum. In the other model, designed to produce (much) larger amounts of radioactive ejecta, we require a highly asymmetric binary (potentially rapidly spinning). These two prior assumptions notably change the conclusions derived from GW170817 and hence the conclusions derived when combining this with other events. Though the latter of our two scenarios seems in tension with the long-lived galactic BNS population, we note that a rapidly-merging subpopulation could have different population properties, and offers an attractive resolution to the long-term challenge of forming r-process elements. Conversely, if we adopt a conventional galactic BNS origin for 170817 and incorporate all available measurements, we conclude $R_{1.4}\simeq 11.8~{}{\rm km}-12.9~{}{\rm km}$.
@@
For simplicity, in this analysis we have omitted several pertinent areas of investigations. We have notably adopted a single GW waveform model, eliding the effect of waveform modeling systematics; see Appendix C for an analysis using TEOBResumS showing that changing the waveform used for EOS inference has a small effect on baseline conclusions, in contrast to changes in the prior assumptions used. Of course, GW waveform models are only calibrated over a certain range of mass ratio, spins, and tides; for the purposes of illustration, at least one of our calculations (the use of high-mass-ratio binaries) stretches the regime of validity of the NRTidalV2 approximation. That said, our headline result – that NS population priors can strongly impact EOS inference, by changing the interpretation of GW170817 – also follows from another example, using comparable masses and positive spins, well within the calibration regime of the NRTidal method. We therefore anticipate the effect size of NS population priors to remain large even as GW waveform models improve. Also, given substantial modeling systematics, we have also only adopted toy models for the information provided from kilonova observations which limit the possible BNS mass, spin, and EoS. We have taken all existing fiducial galactic pulsar measurements at face value; reanalyses of these observations may point to larger systematics and broader posteriors than applied here [163]. Finally, we have adopted an intentionally simple EoS model family and prior, omitting often-discussed constraints available at low density (e.g., from chiral effective field theory and from state-of-the-art crust models) and at high density (e.g, that high-density EoS should asymptotically attain the perturbative-QCD limit of $c_{s}\rightarrow 1/\sqrt{3}$ approaching from below [164, 165]). Our choice of these simplifications enables us to focus exclusively on the often strong and hidden impact that hidden astrophysical priors produce.
@@
The multifactor iterative inference algorithm employed in this work is available as the HyperPipe component of the RIFT python software package, available at https://pypi.org/project/RIFT/. Instructions for this approach are available at [166]. Appendix D illustrates how the hyperpipeline accumulaties information iteratively, in a manner similar to RIFT.
@@ Appendix A NICER data and Gaussian approximation
The NICER data releases [167, 168, 169, 170] provide weighted posterior samples $(x_{k},w_{k})$ derived from comparing their pulse models and neutron star mass-radius combinations with observed X-ray photons. We fit a conventional two-dimensional normal distribution to these weighted samples. For J0030, which lacks any hard edges introduced by causality constraints in the $(M,R)$ plane, we use the weighted empirical mean $\bar{x}=\sum_{k}w_{k}x_{k}/\sum_{k}w_{k}$ and weighted empirical covariance matrix to infer the parameters of the best-fitting gaussian. More generally, however, we optimize the mean $\mu$ and covariance matrix $\Sigma$ of a standard normal distribution, according to
$\ell(\mu,\Sigma)=\sum_{k}w_{k}\ln p(x_{k}|\mu,\Sigma)~{}~{}.$ (13)
In the limit of many weighted samples constructed by importance sampling, with $w_{k}=L(x_{k})$, the quantity $\ell$ will approach the KL divergence between the (truncated) Gaussian and the full posterior. As a check on our results and their stability, we repeat our analysis with fair draws from the weighted samples, using the expression above with $w_{k}=1$. The best fitting normal distribution in $(R/{\rm km},~{}M/{\rm M_{\odot}})$ has a mean $(13.34537396,~{}1.44968231)$ and a covariance matrix
$\begin{bmatrix}1.52687831&0.18343073\\ 0.18343073&0.02778826\end{bmatrix}~{}~{}~{}.$